# Low Resource Video Super-resolution using Memory and Residual Deformable Convolutions

### v1

@@ 3.9 Family of Architectures
    RCDM-light : The architecture consists of same components as our main architecture with the exception of using a lighter feature extraction layer
@@
    RCDM - Our model using deformable convolutions, wavelet transforms, and memory tensors for spatial and temporal consistency.
@@
    RC2DM- Early fusion variant, which integrates contextual information from previous frames at an earlier stage.
@@
    RCDMdwtState - Wavelet-based memory integration for improved long-term temporal coherence.
@@
    RC2DMdwtState - DWT-based feature enhancement of memory tensor for preserving fine-grained details.
@@
RC2DMdwtState 2.10 262.05 0.9075
RC2DM 2.13 272.57 0.9066
3.10 Performance and Computational Trade-offs
The proposed RCDM model achieves state-of-the-art accuracy while remaining computationally efficient. Unlike transformer-based methods such as VRT and RVRT, which require billions of FLOPs, our approach leverages residual convolutions, deformable motion alignment, and memory-based feature propagation to maintain a balance between accuracy and efficiency.
@@ 4.1 Dataset and Training Setup
We train our model on the REDS dataset [13], which is widely used for video super-resolution and restoration tasks. The dataset consists of 300 training sequences, each containing 100 frames with 720p resolution. To enhance the generalization capability of our model, we perform transfer learning using Vimeo90K [14] and YouHQ datasets. These datasets offer diverse motion characteristics and scene complexities, making our model adaptable to real-world scenarios. For evaluation, we test our model on multiple benchmark datasets: REDS4[13], Vimeo90K-T [14], Vid4[18], UDM10[19], and SPCMS[20].
@@ 4.2 Implementation Details
For feature extraction, we leverage ConvNeXt blocks [11] instead of transformers, ensuring reduced computational complexity while retaining long-range dependencies. Our model integrates deformable convolutions [dai2017deformable] for motion compensation, wavelet-based multi-scale feature extraction [15], and a memory tensor module for preserving temporal coherence across frames.
@@ 4.3 Results
Table 2 presents a comparison of various video super-resolution methods across multiple datasets using the SSIM metric. Our proposed RCDM model achieves state-of-the-art results on multiple datasets while maintaining significantly fewer parameters than transformer-based architectures.
Dataset #Params (M) BI Degradation (Bicubic Downsampling) BD Degradation (Blur-Downsampling) REDS4 (RGB) Vimeo90K-T Vid4 UDM10 Vimeo90K-T Vid4 Bicubic - 0.7292 - - - - - TOFlow [24] - 0.7990 0.9054 - - - - EDVR [2] 20.6 0.8800 0.9489 - - - - BasicVSR++ [chan2022basicvsr++] 7.32 0.9069 0.9500 0.8400 0.9722 0.9550 0.8753 EvTexture[22] 8.90 0.9174 0.9544 0.8909 - - - RVRT [4] 11.06 0.9113 0.9527 0.8462 0.9729 0.9576 0.8810 PSRT Recurrent 13.37 0.9106 0.9536 0.8485 - - - VRT [3] 28.80 0.8795 0.9530 0.8425 0.9737 0.9584 0.9006 RCDM (Ours) 2.30 0.9175 0.9212 0.7766 0.9502 0.9194 0.8541
Table 2: Comparison of VSR models across multiple datasets (Other Method’s metrics are as stated in source)
4.3.1 Key Takeaways
  • State-of-the-Art Performance on REDS4: Our method achieves an SSIM of 0.9175, surpassing BasicVSR++ (0.9069) and RVRT (0.9113).
  • Competitive Performance on Vimeo90K-T and UDM10: While transformer-based methods such as VRT and RVRT have strong results, our approach remains competitive while utilizing fewer parameters and significantly lower computational complexity.
  • Memory Tensor Enhancements: The introduction of memory tensors plays a key role in preserving temporal consistency, significantly improving SSIM scores across multiple datasets.
  • Wavelet-Based Feature Extraction Improves Detail Recovery: The use of wavelet transforms contributes to improved structural detail restoration, resulting in better generalization across datasets such as Vimeo, UDM10, and SPCMS.
  • Low Computational Overhead Compared to Transformers: Transformer-based models such as VRT (30.7M parameters) and RVRT (11.06M parameters) exhibit high computational costs, while our model delivers comparable or better performance at significantly lower resource requirements.
5 Qualitative Results and Analysis
To further demonstrate the effectiveness of our model, we present qualitative comparisons against bicubic upsampling across diverse datasets. Figure 3 showcases the patch-wise results on different scenes adopted from SPCMS[20] dataset
@@ 5 Qualitative Results and Analysis
Figure 3: Patch-wise comparison across different samples of SPMCS dataset. First column: Input LR frame (with overlayed patch), Second: Bicubic Upsampling, Third: Our Model Output.
5.1 Inference from the Results
The qualitative comparisons in Figure 3 highlight the superiority of our model in capturing fine details while maintaining computational efficiency. Below are the key observations:
Preservation of Fine Details: - Our model effectively restores sharp textures and edges, as observed in cactus spikes, car headlights, and building edges. - Unlike bicubic interpolation, which produces blurry and pixelated results, our method reconstructs clear and natural textures.
High-Quality Super-Resolution Without Transformers: - Unlike transformer-based approaches, which introduce global self-attention, our model preserves local structures using ConvNeXt blocks and deformable convolutions. - Transformers are beneficial for high-level vision tasks (e.g., segmentation) but are unnecessary for low-level vision tasks like super-resolution, where local feature extraction is more effective.
Lower Computational Complexity Without Sacrificing Quality: - By leveraging memory mechanisms and efficient convolutions, our approach achieves comparable or superior quality with significantly lower computational overhead than transformer-based methods. - This makes our model highly practical for real-time applications, including edge devices and video processing.
Robust Performance Across Textures and Patterns: - Our model generalizes well to natural textures, urban environments, and structured patterns, delivering consistent performance across diverse datasets.
Our results confirm that global self-attention is not always necessary for video super-resolution, and efficient local convolutions can achieve high-quality reconstruction with lower computational costs. Our model is a highly practical alternative to transformer-based VSR approaches particulary for edge devices.
@@ 6 Ablation Study
To systematically evaluate the contributions of different components in our proposed architecture, we conducted an extensive ablation study. This study assesses the impact of architectural modifications on model complexity (measured in parameters) and performance (measured using the Structural Similarity Index Measure, SSIM). The findings are summarized in Table LABEL:tab:ablation_study.
6.1 Key Observations
1. Effect of Architectural Modifications
Our baseline ResConvDefMem model (2.30M parameters) achieves an SSIM of 0.9175, outperforming several larger models while maintaining a low parameter count.
Replacing ConvNeXt with Swin Transformer for upscaling the frame (2.69M parameters) results in a negligible change in SSIM (0.9170), confirming that transformer-based architectures are not essential for video super-resolution, which is inherently a local feature enhancement task.
2. Impact of Multi-Scale and Wavelet-Based Processing
The introduction of wavelet-based multi-scale feature extraction in RCDMdwtState leads to an improvement in SSIM (0.9133) compared to its standard counterpart (0.9056 SSIM). This indicates that frequency-domain information enhances fine-detail preservation in super-resolution tasks.
3. Model Complexity vs. Performance Trade-off
Reducing convolutional layers (lighter models) leads to slight degradation in SSIM, as seen in the RCDm-light (1.86M parameters, SSIM 0.9041).
However, training on specific datasets such as VimeoBI and VimeoBD achieves strong results with minimal parameter finetuning (0.60M parameters, SSIM 0.9212), reinforcing that our model has high transferability.
These findings validate our approach of designing a memory-enhanced, wavelet-based deformable convolutional model, which attains superior performance at a fraction of the computational cost of transformer-based models.
Method Model Parameters (M) SSIM
RCDM (Original) 2.30 0.9175
ResSwinMem 2.69 0.9170
RCDM Light Recurrent 1.27 0.7772
RCDM Recurrent 2.23 0.8056
Swin Recurrent 2.48 0.8080
RCDM (without Memory) 2.30 0.9051
RCDM (without Wavelet) 2.28 0.9062
YouHQ Dataset 0.60 0.6768
VimeoBD Dataset 0.60 0.9192
VimeoBI Dataset 0.60 0.9212
Table 3: Comparison of Performance of Model and Dataset Variations
We also performed experiments in YCbCr space instead of RGB space to compare model performance. However, for our dataset of interest (REDs), model gave better performance in the RGB space.
@@ 7.1 Conclusion
Our experiments demonstrate that RCDM achieves nearly state-of-the-art performance on the REDS4 dataset with an SSIM of 0.9175, surpassing existing lightweight models and even competing with several high-resource transformer-based architectures such as VRT and RVRT.
Furthermore, our qualitative and quantitative results validate the generalizability across diverse datasets. By prioritizing convolutional efficiency over transformer-based global attention, our model offers a practical solution for real-time applications such as video streaming, surveillance, and edge computing.
7.2 Future Work
Despite achieving competitive SSIM scores, we observed that in highly dynamic scenes with fast motion, our model sometimes struggles to recover sharp visual details. While memory-based feature propagation improves temporal consistency, further enhancements are needed to ensure finer detail preservation and structure reconstruction in challenging scenarios.
For future research, we aim to explore some of the following. Perceptual Loss Optimization: Experimenting with perceptual loss functions such as LPIPS and adversarial losses to enhance the subjective visual quality of reconstructed frames. Also, in updating the current memory tensor to assist in further motion compensation.
By addressing these areas, we aim to further refine RCDM to maximize both perceptual quality and computational efficiency, making it an optimal solution for next-generation video enhancement applications in real-world environments.

### v2

@@ 3.7 Family of Architectures
    RCDM: This is our main model that uses deformable convolutions, wavelet transforms, and memory tensors for spatial and temporal consistency.
@@
    RCDM-light: This architecture consists of same components as our main architecture with the exception of using a lighter feature extraction layer.
@@
    RC2DM: This variant uses early fusion to integrate contextual information from previous frames at an earlier stage.
@@
    RCDMdwtState: This variant uses wavelet-based memory integration for improved long-term temporal coherence.
@@
    RC2DMdwtState: This variant uses DWT-based feature enhancement of memory tensor for preserving fine-grained details.
@@
3.8 Performance and Computation Trade-offs
The proposed RCDM model achieves high accuracy while remaining computationally efficient. Our approach leverages residual convolutions, deformable motion alignment, and memory-based feature propagation to maintain a balance between accuracy and efficiency.
@@ 4.1 Dataset and Training Setup
We trained our model on the REalistic and Dynamic Scenes dataset (REDS) dataset [20], which is widely used for video super-resolution and restoration tasks. The dataset consists of 300 training sequences, each containing 100 frames with 720p resolution. To enhance the generalization capability of our model, we perform transfer learning on Vimeo90K [27] and YouHQ datasets. These datasets offer diverse motion characteristics and scene complexities, making our model adaptable to real-world scenarios. For evaluation, we test our model on multiple benchmark datasets: REDS4[20], Vimeo90K-T [27], Vid4[15], UDM10[28], and SPCMS[24]. We compare structural similary index measure (SSIM) metric across the experiments as PSNR tends to prioritize global intensity over local detail matching.
@@ 4.2 Implementation Details
For feature extraction, we leverage ConvNeXt blocks [18] instead of transformers, ensuring reduced computational complexity while retaining long-range dependencies. Our model integrates deformable convolutions [7] for motion compensation, wavelet-based multi-scale feature extraction [10], and a memory tensor module for preserving temporal coherence across frames.
@@ 4.3 Results
Table 2 presents a comparison of various video super-resolution methods across multiple datasets using the SSIM metric. Our proposed RCDM model achieves results comparable to the state-of-the-art on multiple datasets while using far fewer parameters (Table 2.
Table 2: Comparison of VSR models across multiple datasets (Other method’s metrics are taken from their respective papers.)
Dataset #Params (M) BI Degradation (Bicubic Downsampling) BD Degradation (Blur-Downsampling) REDS4 (RGB) Vimeo90K-T Vid4 UDM10 Vimeo90K-T Vid4 Bicubic - 0.7292 - - - - - TOFlow [26] - 0.7990 0.9054 - - - - EDVR [25] 20.6 0.8800 0.9489 - - - - BasicVSR++ [3] 7.32 0.9069 0.9500 0.8400 0.9722 0.9550 0.8753 EvTexture[12] 8.90 0.9174 0.9544 0.8909 - - - RVRT [14] 11.06 0.9113 0.9527 0.8462 0.9729 0.9576 0.8810 PSRT Recurrent[22] 13.37 0.9106 0.9536 0.8485 - - - VRT [13] 28.80 0.8795 0.9530 0.8425 0.9737 0.9584 0.9006 RCDM (Ours) 2.30 0.9175 0.9212 0.7766 0.9502 0.9194 0.8541
Our method demonstrates impressive performance on the REDS4 dataset, surpassing BasicVSR++ (0.9069) and RVRT (0.9113). Additionally, it remains competitive on Vimeo90K-T and UDM10, despite transformer-based models like VRT and RVRT showing strong results. Notably, our approach achieves this with fewer parameters and inference compute (Figure 1).
To further demonstrate the effectiveness of our model, we present qualitative comparisons against bicubic upsampling across diverse datasets. Figure 3 presents the patch results in different scenes adopted from the SPCMS[24] dataset, where fit can be seen that fine details are recreated with sharpness and without artifacts.
@@
Figure 3: Patch-wise comparison across different samples of the SPMCS dataset. First column: input LR frame (with overlaid patch), Second: bicubic upsampling, Third: our model’s output.
[…]
Bicubic Refer to caption
RVRT Refer to caption
BasicVSR++ Refer to caption
Ours Refer to caption
@@
Figure 4: Patch-wise comparison across different samples. Each row represents a different case: (1) Low-Resolution input with an enlarged overlaid patch, (2) Bicubic upsampling, (3) RVRT output, (4) BasicVSR++ output, (5) Our output.
4.4 Analysis of Results
Our model demonstrates superior detail preservation compared to Bicubic Upsampling, RVRT, and BasicVSR++. The fine structures in architectural patterns, edges, and textures are sharply reconstructed, while other methods exhibit blurring or texture loss. The inclusion of memory tensors and wavelet-based feature extraction ensures that the generated frames maintain both temporal consistency and high-frequency details, outperforming transformer-based approaches that often introduce artifacts or over-smoothing. These results confirm the effectiveness of our method in real-world scenarios where fine-grained structures and text clarity are critical.
Our model performs well on the the REDS4 dataset as it is a long sequence along with dynamic motion. Table 3 shows the nature of the datasets under consideration. Our model performs with dataset with high motion similar to REDS
Table 3: Comparison of Motion Complexity Across Datasets
Dataset Motion Complexity Frame Rate Key Challenges
REDS4[20] Very High 24 fps - Fast-moving objects, occlusions.
- Large scene variations.
UDM10[28] Moderate - High 30 fps - Real-world compression artifacts.
- Camera jitter, lossy motion.
Vimeo-90K[27] Moderate  30 fps - Well-curated natural motion.
- Fewer occlusions, predictable motion.
Vid4[15] Low 24 fps - Mostly slow camera panning.
- Minimal motion complexity.
The qualitative comparisons in Figure 4 highlight the superiority of our model in capturing fine details while maintaining computational efficiency. Below are the key observations: Preservation of Fine Details: - Our model effectively restores sharp textures and edges, as observed in cactus spikes, car headlights, and building edges. - Unlike bicubic interpolation, which produces blurry and pixelated results, our method reconstructs clear and natural textures. High-Quality Super-Resolution Without Transformers: - Unlike transformer-based approaches, which introduce global self-attention, our model preserves local structures using ConvNeXt blocks and deformable convolutions. - Transformers are beneficial for high-level vision tasks (e.g., segmentation) but are unnecessary for low-level vision tasks like super-resolution, where local feature extraction is more effective. Lower Computational Complexity Without Sacrificing Quality: - By leveraging memory mechanisms and efficient convolutions, our approach achieves comparable or superior quality with significantly lower computational overhead than transformer-based methods. - This makes our model highly practical for real-time applications, including edge devices and video processing. Robust Performance Across Textures and Patterns: - Our model generalizes well to natural textures, urban environments, and structured patterns, delivering consistent performance across diverse datasets.
Our results confirm that global self-attention is not always necessary for video super-resolution, and carefully crafted local architectures can achieve high-quality reconstruction with lower computational costs. Our model is a highly practical alternative to transformer-based VSR approaches particularly for edge devices.
@@ 5 Ablation Study
To systematically evaluate the contributions of different components in our proposed architecture, we conducted an extensive ablation study. This study assesses the impact of architectural modifications on model complexity (measured in parameters) and reconstruction accuracy (measured in SSIM).
(a) SSIM Improvement using Memory Mechanism
(b) SSIM Improvement using Wavelet-Based Feature Extraction
Figure 5: Ablation Study: (a) Impact of memory mechanism and (b) Impact of wavelet-based feature extraction on SSIM performance.
Impact of Memory and Wavelet Transform: The ablation study evaluates the effectiveness of two key components in our Residual ConvNeXt Deformable Memory (RCDM) model: memory mechanism and wavelet-based feature extraction. Figure 5 presents the SSIM ratio for both components (using an architecture without these components as baseline), showing their contribution to enhancing video super-resolution quality on a subset of REDS4 dataset.
Figure 5a shows that memory propagation significantly improves temporal consistency, especially in later frames as more evidence to resolve patches is accumulated by the memory. Thus, the memory mechanism helps retain long-range dependencies, reducing flickering and enhancing detail retention in dynamic scenes.
Figure 5b shows that the wavelet decomposition improves high-frequency detail preservation, leading to better texture reconstruction. Unlike standard convolutions, wavelet transforms effectively separate structural information across different frequency bands, enabling finer detail recovery. This results in higher SSIM ratios, especially in complex background regions. Frame-to-frame consistency benefits from memory propagation, particularly in high-motion sequences. Wavelets enhance texture reconstruction: By leveraging frequency decomposition, wavelet-based models preserve edges and fine details more effectively.
In fact, the memory and wavelet components complement each other. Memory helps smooth temporal transitions, while wavelets refine structural integrity. Their combined use yields optimal results in video super-resolution.
We also performed experiments in YCbCr space instead of RGB space to compare model performance. However, for our dataset of interest (REDS4), the model gave better performance in the RGB space.
@@ 6 Conclusion and Future Work
Our experiments demonstrate that RCDM achieves reconstruction quality comparable to the state-of-the-art on several datasets and surpassing SOTA on the REDS4 dataset with an SSIM of 0.9175 when comapred to both lightweight VSR models and high-resource transformer-based architectures such as VRT and RVRT. Furthermore, our qualitative and quantitative results validate the generalizability across diverse datasets. By prioritizing convolutional efficiency over transformer-based global attention, our model offers a practical solution for real-time applications such as video streaming, surveillance, and edge computing.
While memory-based feature propagation improves temporal consistency, further enhancements are needed to ensure finer detail preservation and structure reconstruction in challenging scenarios. We can refine RCDM, perhaps using locally limited self-attention (e.g. that in SWIN transformer [17], to maximize both perceptual quality and computational efficiency, making it an optimal solution for next-generation video enhancement applications.